# Large Learning Rates Simultaneously Achieve Robustness to Spurious Correlations and Compressibility

### v1

@@ D.1 Proof of Proposition 1
as $k\to\infty$. As $k\to\infty$, $-\log(\pi_{y})\approx-\log(1-e^{-k\alpha})\approx e^{-k\alpha}$. Assume (without loss of generality) that the margins $\beta$ and $\alpha$ are shared by all bias-conflicting and bias-aligned samples in the minibatch. Then, as $k\to\infty$,
@@ D.2 Additional Theoretical Results
$f_{\mathbf{w}}$ is called a bias-decomposable network if $f_{\mathbf{w}}(\mathbf{x})=f_{\mathbf{c}}(\mathbf{x})+f_{\mathbf{s}}(\mathbf{x})+f_{\mathbf{n}}(\mathbf{x})$. Here, $f_{\mathbf{c}}$ is the core feature subnetwork, $f_{\mathbf{c}}(\mathbf{x})[y]-f_{\mathbf{c}}(\mathbf{x})[b]\geq 0$ with equality iff $y=b$. $f_{\mathbf{c}}(\mathbf{x})$ is assumed to have converged to a stable decision making rule, i.e. $\nabla_{\mathbf{c}}(f_{\mathbf{c}}(\mathbf{x})[y]-f_{\mathbf{c}}(\mathbf{x})[b])\approx 0$. In contrast, $f_{\mathbf{s}}$ is the spurious feature subnetwork, $f_{\mathbf{s}}(\mathbf{x})[y]-f_{\mathbf{s}}(\mathbf{x})[b]\leq 0$ with equality iff $y=b$.
Note that although this decomposition is inevitably a simplification, it is a reasonable one in light of previous findings that demonstrate such modularity [31], and similar decompositions have been utilized in related research previously [57]
In the following proposition, we investigate how the gradient norms for core and spurious subnetworks scale based on this definition and the results in the main paper.
@@
[…]
@@
The proof follows from the gradient of the cross-entropy loss with respect to the parameters $\textbf{w}_{c}$ and $\textbf{w}_{s}$. The softmax probability $\pi_{j}$ for a class $j$ is given by:
@@
The gradient of the loss $\ell$ can be expressed as:
@@ D.3 Value Ranges for Figures
Given that our figures depict multiple variables at the same time, and the results are normalized according to experiments to illuminate the patterns that LR and other interventions create, we present the min and max values the independent and dependent variables take in Tab. 1 and Tab. 2.
@@ D.4 Additional Experiment Results
[…]
@@
Figure 24: (Top) Effects of learning rate on OOD performance (unbiased test acc.), network prunability, and representation properties when trained for $100\%$ training accuracy and $<0.00001$ training loss. (Bottom) Test accuracy does not meaningfully change beyond convergence (vertical lines correspond to the point where 100% was reached).
[…]
@@
We confirm our findings on robustness to SCs and comp. extend to modern and standard training setups. Fig. 28 (top, left) shows that core benefits persist with ResNet50 (Colored MNIST) using the PSGD (Kron) optimizer and a WSD LR schedule. Fig. 28 (top, center) shows that our key findings also hold under a standard CIFAR-100 setup with AdamW, cosine annealing, weight decay, and validation set based model selection, addressing concerns about reliance on constant LR SGD. The effects of LR were even more dramatic in both cases, e.g. in the former, the range between low vs. high LRs was $\sim$22- $\sim$91compared to the original $\sim$55- $\sim$86%. Fig. 28 (top, right) also shows the performance of the final model after LR reduction by iteration $i$: the effects of LR are almost completely integrated by step 1000.
Spurious feature utilization of a single neuron
Fig. 28 (bottom) plots the spurious feature utilization (SFU; computed via attribution methods) of a single neuron under high LRs. The neuron is effectively “reset” with large updates as its SFU increases. Our revision will include a class and bias specific analyses of neuron gradients.
Figure 28: Additional experiments and analyses.
@@ E Additional Details and Results Regarding Neural Network Attribution
E.1 Attribution Methods in Neural Networks
Attribution methods are a prominent methodology within explainable artificial intelligence (XAI) [47]. Given a fixed predictor, e.g. a trained neural network, such methods aim to measure the sensitivity of the neural networks’ internal representations or output to changes in the input or internal representations. The most straightforward notion of attribution involves the relationship between input and output (called primary or input attribution). For example, in an image classifier, this corresponds to quantifying the sensitivity of the prediction with respect to every pixel and channel. Various methods with their unique advantages and disadvantages have been proposed in the literature; see [1].
@@ E.1 Attribution Methods in Neural Networks
[…]
@@
We train a ResNet18 model using a low vs. high LR with 10 different seeds on CIFAR-10 and CIFAR-100 datasets. Then, we extract those samples in the test set which have been correctly predicted by >.$>.75$ of the high LR models and <.$<.25$ of low LR ones. Then, we investigate the attribution maps of low vs. high LR models in Fig. 8.
E.2 Additional Attribution Visualizations
We provide additional visualization of attributions for our experiments in the main paper; for Colored MNIST (Fig. 29), MNIST-CIFAR (Fig. 30), Double MNIST (Fig. 31), CelebA (Fig. 32), CIFAR-10 (Fig. 33), and CIFAR-100 (Fig. 34) datasets. Notice that as in the main paper, low learning rate (LR) models are more likely to focus on spurious features compared to high LR models.

### v2

@@ D.1 Proof of Proposition 1
as $k\to\infty$. As $k\to\infty$, $\log(\pi_{y})\approx-\log(1-e^{-k\alpha})\approx e^{-k\alpha}$. Assume (without loss of generality) that the margins $\beta$ and $\alpha$ are shared by all bias-conflicting and bias-aligned samples in the minibatch. Then, as $k\to\infty$,
@@ D.2 Additional Theoretical Results
$f_{\mathbf{w}}$ is called a bias-decomposable network if $f_{\mathbf{w}}(\mathbf{x})=f_{\mathbf{c}}(\mathbf{x})+f_{\mathbf{s}}(\mathbf{x})+f_{\mathbf{r}}(\mathbf{x})$. Here, $f_{\mathbf{c}}$ is the core feature subnetwork, $f_{\mathbf{c}}(\mathbf{x})[y]-f_{\mathbf{c}}(\mathbf{x})[b]\geq 0$ with equality iff $y=b$. $f_{\mathbf{c}}(\mathbf{x})$ is assumed to have converged to a stable decision making rule, i.e. $\nabla_{\mathbf{c}}(f_{\mathbf{c}}(\mathbf{x})[y]-f_{\mathbf{c}}(\mathbf{x})[b])=0$. In contrast, $f_{\mathbf{s}}$ is the spurious feature subnetwork, $f_{\mathbf{s}}(\mathbf{x})[y]-f_{\mathbf{s}}(\mathbf{x})[b]\leq 0$ with equality iff $y=b$. $f_{\mathbf{r}}$ is the remainder subnetwork that does not conform to the behavior described for $f_{\mathbf{c}},f_{\mathbf{s}}$.
Before discussing the motivation for this idealization, in the following proposition, we investigate how the gradient norms for core and spurious subnetworks scale based on this definition and the results in the main paper.
@@
for some $\alpha>0$, where $\|\cdot\|$ stands for Frobenius norm.
@@
The softmax probability $\pi_{j}$ for a class $j$ is given by:
@@
With some abuse of notation, the gradient of the loss $\ell$ can be expressed as:
@@
Note that although the decomposition in question is an idealization, it follows in line of previous work that demonstrate such modularity [32], and similar decompositions have been utilized in related research previously [60]. However, note that our assumption regarding the stability of the core vs. spurious subnetworks differ from that of [60], who assume a Bayes-optimal, stable spurious subnetwork. Our assumption is motivated by our empirical observations. In Fig. 18 (top), we examine two ResNet18 models trained on Colored MNIST dataset, and investigate their learned representations, based on the pooled outputs of the last layer convolution filters. We compute the CSI - BSI values for each filter/neuron at different points in the training. Assume we categorize neurons for which CSI >> BSI as “class-dominant”, and “bias-dominant” otherwise. We then ask, “At iteration $t$, what percentage of the neurons that were class dominant remained class-dominant by the end of training, and what percent of the neurons that were bias-dominant remained so?” As the results show, for high learning rate ($\eta=0.1$ vs. $\eta=0.001$), after iteration $\sim 10$ the overwhelming majority (not infrequently 100%) of the neurons who were class-dominant remained so. This is not true for the bias-dominant neurons. Note that this is not explained solely by the final ratio of class-dominant neurons: 55.4% vs 85.9% for the two models respectively, which falls short of explaining the behavior observed. To provide a more microscopic examination of this, in Fig. 18 (bottom) we examine how neuron-specific gradients impact spurious feature utilization (computed through feature attribution) within that neuron, from a high LR ($\eta=0.1$) experiment within our FCN + Colored MNIST setting. The updates clearly show that the increasing spurious feature utilization is “reset” by a large gradient update, driven presumably by mispredictions from bias-conflicting samples. We consider further examination of these phenomena as an important future research direction.
Figure 18: (Top) Ratio of class-dominant vs. bias-dominant neurons that survived as such by the end of training. (Bottom) Examining the impact of large gradient updates on feature attributions.
@@ D.3 Value Ranges for Figures
Given that our figures depict multiple variables at the same time, and the results are normalized according to experiments to illuminate the patterns that LR and other interventions create, we present the min. and max. values the independent and dependent variables take in Tab. 1 and Tab. 2.
@@ D.4 Additional Experiment Results
To show that our results are not limited to training using SGD with a constant LR, we present qualitatively identical experiment results in Fig. 23 using the Colored MNIST dataset and ResNet18 model, where the initial learning rate ($x$-axis) is multiplied by 0.1 after 1000th iteration. Similar results using Adam optimization algorithm are presented in Fig. 24. Additionally, using the same setting, in Fig. 25 we show that training models for longer according to an additional criterion (CE loss <1e5<1e-5) produces qualitatively identical results as test accuracy changes very little beyond convergence for both low and high LR models. Finally, Fig. 26 demonstrates that alternative choices to characterize parameter and representation compressibility, such as $(q,\kappa)\textrm{-Compressibility}$, sparsity, and the recently proposed PQ-Index [13] produce qualitatively identical results.
@@
Figure 25: (Top) Effects of learning rate on OOD performance (unbiased test acc.), network prunability, and representation properties when trained for $100\%$ training accuracy and <0.00001<0.00001 training loss. (Bottom) Test accuracy does not meaningfully change beyond convergence (vertical lines correspond to the point where 100% was reached).
Figure 26: Utilizing alternative notions of parameter and representation compressibility such as prunability, $(q,\kappa)\textrm{-Compressibility}$ (with $q=2,\kappa=0.1$), sparsity, and the recently proposed PQ-Index (with $p=2,q=1$).
@@
We confirm our findings on robustness to SCs and comp. extend to modern and standard training setups. Fig. 29 (left) shows that core benefits persist with ResNet50 (Colored MNIST) using the PSGD (Kron) optimizer and a WSD LR schedule. Fig. 29 (right) shows that our key findings also hold under a standard CIFAR-100 setup with AdamW, cosine annealing, weight decay, and validation set based model selection, addressing concerns about reliance on constant LR SGD.
Fig. 30 (left) compares ResNet18 models trained on CIFAR-10 with a starting LR of $0.1$, which is decreased to $0.001$ by step $S$. We plot the eventual unbiased test performance of a model as function of $S$, with the performance of a model trained by a constant LR of $0.1$ vs. $0.001$ depicted as horizontal lines for reference. The results show that the effects of LR are almost completely integrated by $S=1000$. Fig. 30 (right) shows that creating a “model soup” [81] is another way of obtaining robustness vs. compressibility disentanglement.
Figure 29: Experiments with alternative optimizers, schedulers, and convergence criteria.
Figure 30: (Left) Unbiased test set performance as a function of LR reduction. (Right) Disentangling compressibility and robustness through a “model soup” (the rightmost model).
@@ E Additional Details and Results Regarding Neural Network Attribution
$\mathrm{IG}_{i}(\mathbf{x}):=(x_{i}-x^{\prime}_{i})\times\int_{\alpha=0}^{1}\frac{\partial f(\mathbf{x}^{\prime}+\alpha(\mathbf{x}-\mathbf{x}^{\prime}))}{\partial x_{i}}d\alpha.$
@@
We train a ResNet18 model using a low vs. high LR with 10 different seeds on CIFAR-10 and CIFAR-100 datasets. Then, we extract those samples in the test set which have been correctly predicted by >.75>.75 of the high LR models and <.25<.25 of low LR ones. Then, we investigate the attribution maps of low vs. high LR models in Fig. 8.
E.1 Additional Attribution Visualizations
We provide additional visualization of attributions for our experiments in the main paper; for Colored MNIST (Fig. 31), MNIST-CIFAR (Fig. 32), Double MNIST (Fig. 33), CelebA (Fig. 34), CIFAR-10 (Fig. 35), and CIFAR-100 (Fig. 36) datasets. Notice that as in the main paper, low LR models are more likely to focus on spurious features compared to high LR models.